OpenStudy (anonymous):

A driver travels 372 miles in 6 hours. At that rate, how far will the driver travel in 15 hours?

OpenStudy (anonymous):

372/6 = x/15 (372 miles to 6 hrs = x miles to 15 hrs) (6)(x) = (372)(15) 6x = 5580 -- divide by 6 to get x by itself) x = 930

sammixboo (sammixboo):

You have choices?

OpenStudy (texaschic101):

LegoMyEgo is correct...put it in a proportion

OpenStudy (gudden):

speed = distance upon time and here speed comes out to be 372/6 = 62 mi/hour and distance = speed * time therefore distance thus travelled in 15 hours = 62 * 15 = 930 mi

sammixboo (sammixboo):

Ok yes 930 is correct, and I got this the way that @gudden did it.

OpenStudy (anonymous):

Sorry I will try to explain it to your more next time.

OpenStudy (gudden):

@LegoMyEgo .. you just didn't mention that what were u going to do the next... otherwise ur answer as correct

OpenStudy (texaschic101):

I would have done it the way LegoMyEgo did it

OpenStudy (texaschic101):

either way is correct

sammixboo (sammixboo):

Well everyone has there own techniques. :) And yeah they are both correct.

OpenStudy (texaschic101):

so true :)

OpenStudy (gudden):

@texaschic101 .. i would have also done that way... but the i realized the fact that the asker may be new so he/she couldnot understand the short cut way...

OpenStudy (texaschic101):

thats true too.